Circular No. 8043

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003C IN UGC 439 Further to IAUC 8037, T. Puckett and L. Cox report the discovery of

a supernova (mag 17.0) on an unfiltered CCD frame (limiting mag 20.5) taken with the 0.60-m automated patrol telescope on Jan. 6.02 UT (and confirmed on CCD frames taken by J. Newton, Portal, AZ, on Jan. 7.04). SN 2003C is located at α = 0^{h}41^{m}20^{s}.91, δ = −1^{o}42′54′′.1 (equinox 2000.0), which is 8′′.8 west and 2′′.8 north of the center of UGC 439. The new object was not present on images taken by Puckett on 2002 Aug. 4 and Sept. 6 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken 1982–1992.

A. V. Filippenko and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Jan. 7 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2003C is of type II, shortly after the explosion. Low-contrast lines of the hydrogen Balmer series and He I 587.6-nm exhibit P-Cyg profiles superposed on a very blue continuum.

SUPERNOVA 2003D IN MCG –01-25-9 Puckett and A. Langoussis also report the discovery of a supernova

(mag 17.5) on an unfiltered CCD frame taken (as above) on Jan. 6.29 UT (confirmed on Jan. 7.16). SN 2003D is located at α = 9^{h}38^{m}53^{s}.52, δ = −4^{o}51′05′′.1 (equinox 2000.0), which is 1′′.6 east and 10′′.2 south of the center of MCG –01-25-9. The new object was not present on images taken by Puckett on 2002 Feb. 14 (limiting mag ∼ 20.5) or on Palomar Sky Survey plates taken in 1983 and 1985 (limiting mag ∼ 20.0–21.0).

Filippenko and Chornock add: “Keck spectra obtained as above reveal that the object is a peculiar type-Ia supernova not long after maximum brightness. It is spectroscopically similar to the subluminous SN 1991bg (e.g., Filippenko et al. 1992, A.J. 104, 1543), but with some distinct differences. For example, its Ti II trough around 420 nm is not as strong as that of SN 1991bg, while the 460-nm emission feature is much stronger. The expansion velocities of the two objects are very similar, ∼ 10000 km/s.”

SUPERNOVA 2003B NEAR NGC 1097 Precise CCD position for SN 2003B (cf. IAUC 8042) by K. Kadota,

Ageo, Saitama, Japan (via S. Nakano), from a CCD image taken on Jan. 7.465 UT (when V = 15.6): α = 2^{h}46^{m}13^{s}.78, δ = −30^{o}13′45′′.1 (equinox 2000.0), or 68′′ west and 163′′ north of the center of NGC 1097.

2003 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green